# CoViS-Net: A Cooperative Visual Spatial Foundation Model for Multi-Robot Applications

### v1

@@ Appendix B Real-World Dataset
Figure 12: We collect the real-world dataset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$ from five unique scenes in an indoor office building. The scenes show different challenges, from cluttered environments over scenes with high ceilings to sunny floors. We show the distribution of positions for each scene (second row) and the distribution of relative angles between nodes (third row).
Appendix C Results
C-A Simulation
We visualize four additional samples from the simulation dataset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$ in Fig. 14, Fig. 15, Fig. 16, Fig. 17, Fig. 18 and Fig. 19.
Figure 14: Sample A from $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$.
Figure 15: Sample B from $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$.
Figure 16: Sample C from $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$.
Figure 17: Sample D from $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$.
Figure 18: Sample E from $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$.
Figure 19: Sample F from $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$.
C-B Real-World
We visualize two more samples from the real-world dataset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$ in Fig. 20, Fig. 21, Fig. 22, Fig. 23.
Figure 20: Sample A from $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$.
Figure 21: Sample B from $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$.
Figure 22: Sample C from $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$.
Figure 23: Sample D from $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$.

### v2

@@ B.4 Real-World Pose Control
Table 6: Model runtime evaluation for $S=128$ and $F=24$.
Type Mode Dev. $f_{\textrm{enc}}$ $f_{\textrm{pose}}$
FP16 TRT GPU 19.22 ms $\pm$0.92 ms 4.15 ms $\pm$0.49 ms
FP16 JIT GPU 36.79 ms $\pm$1.33 ms 8.04 ms $\pm$0.73 ms
FP32 JIT GPU 99.32 ms $\pm$1.56 ms 13.57 ms $\pm$1.10 ms
FP32 JIT CPU 952.04 ms $\pm$116.72 ms 248.41 ms $\pm$45.98 ms
Figure 20: We evaluate the tracking performance of our model and uncertainty-aware controller on two additional reference trajectories, with two follower robots (in blue and orange), positioned left and right of the leader robot (in green). For each trajectory, we report the tracking performance over time for position and rotation, as well as the predicted uncertainties. The top trajectory (1) is a Figure Eight with the leader always facing in the same direction. The bottom trajectory (2) is a rounded rectangle. We show the trajectory for 120 s and the tracking error for the first 60 s. The tracking performance is consistent across multiple runs.
Sec. B.4 reports quantitative evaluations for the same results, which are in line with the results reported in Tab. 1. Note that the velocities reported are average velocities of the follower robots. The velocities are consistent for both Figure 8 trajectories, but differ for the rectangle trajectory, as the inner robot moves more slowly than the outer robot.
Table 7: We report the mean and absolute tracking error for both leaders for all three trajectories, as well as average velocities. The errors are consistent with the results reported in Tab. 1.
Trajectory Robot Mean Abs. Median Vel
Figure 8 dynamic A $38$ cm, $5.6^{\circ}$ $38$ cm, $4.8^{\circ}$ 0.59 m/s
B $28$ cm, $5.1^{\circ}$ $25$ cm, $4.7^{\circ}$ 0.58 m/s
Figure 8 static A $51$ cm, $5.2^{\circ}$ $48$ cm, $5.3^{\circ}$ 0.60 m/s
B $28$ cm, $5.6^{\circ}$ $26$ cm, $5.4^{\circ}$ 0.61 m/s
Rectangle dynanic A $41$ cm, $4.4^{\circ}$ $42$ cm, $4.4^{\circ}$ 0.32 m/s
B $29$ cm, $5.1^{\circ}$ $27$ cm, $5.1^{\circ}$ 0.81 m/s